OpenStudy (anonymous):

Simplify Squareroot of 36

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

6

OpenStudy (angela210793):

=+-6

OpenStudy (anonymous):

no Like not solve. Simplify

OpenStudy (anonymous):

Oh

OpenStudy (amistre64):

to simplify is to put in simplest form; +-6 IS simplest form

OpenStudy (anonymous):

Its +6 or -6

OpenStudy (angela210793):

how can u simplify??? tht's it....+-6

OpenStudy (anonymous):

\[\sqrt{36} = \sqrt{6 x 6} = \sqrt{6^{2}} = 6\]

OpenStudy (anonymous):

So final answer is +6 or -6

OpenStudy (anonymous):

ok thanks Harkirat that helped

OpenStudy (anonymous):

U r welcome....